# Finite Element and Machine Learning Modeling of Autogenous Self-Healing Concrete

### v1

@@ Abstract
A time-dependent modeling framework for autogenous self-healing concrete that couples moisture diffusion with damage evolution was developed. Water transport follows Fick’s second law with a damage-dependent diffusivity obtained by power-law interpolation between intact concrete and crack space. Healing reduces damage in proportion to local moisture and a smoothed cement availability field computed via a Helmholtz filter. Two finite element variants were implemented in FEniCSx over time horizons up to $5\times 10^{6}$ seconds: a Crack Diffusion Model (CDM) with standard diffusion and a Crack Membrane Model (CMM) that gates cross-crack transport until a critical moisture threshold is reached. Key control parameters are the initial crack orientation and size, the diffusion coefficients of intact and cracked concrete, the healing rate constant, and the cement availability smoothing parameter. Simulations on a unit square domain show that healing time varies non-monotonically with crack orientation, peaking near $45^{\circ}$ and $135^{\circ}$ and minimizing near $90^{\circ}$, consistent with diffusion distance to crack endpoints dominating the process. The dependence on crack width reverses with material parameters: healing time increases when $D_{\text{cracked}}<D_{\text{intact}}$ and decreases when $D_{\text{cracked}}>D_{\text{intact}}$. The CMM reproduces staged moisture penetration with delayed gate activation but lengthens total healing time, whereas the CDM is efficient for parametric sweeps. Machine learning classifiers trained on one million simulation samples predict binary healing outcomes $H(\sigma,\gamma,t)$ (healed or not) with high accuracy (up to 0.998 for neural networks). Although experimental calibration is still required, the framework provides a versatile tool for guiding laboratory studies and implementations of self-healing concrete.
@@ 1 Introduction
Concrete is the second most widely used material globally, surpassed only by water, with over 30 billion metric tons utilized worldwide annually [1]. This critical construction material has been employed for over 8,000 years, dating back to concrete-like structures built by Nabataea traders in modern-day Syria [2], and continues to be essential for housing, transportation, and water infrastructure today. The global concrete market is projected to reach \$821.6 billion by 2026 [3], driven by growing investment in construction projects worldwide.
@@
The infrastructure crisis compounds these environmental challenges. Most countries completed their major infrastructure projects during the late 1900s, including the US Interstate Highway System, European motorways, Japan’s Shinkansen, and China’s expressway networks. Over half a century later, these aging structures face increasing failure risks, with concrete failure probability approaching 70% after 50 years and becoming nearly inevitable by the century’s end [7]. Currently, over 20,000 concrete bridges in the United States are classified as structurally deficient [8], requiring frequent maintenance and posing safety risks. The global cost of corrosion in reinforced concrete infrastructure is estimated at \$2.5 trillion annually, representing approximately 3.4% of global GDP [9].
Self-healing concrete emerges as a promising solution to address these challenges. This innovative material repairs its cracks automatically without human intervention, typically through chemical or biological processes [10]. Self-healing concrete offers reduced maintenance requirements over infrastructure lifetime, diminished social costs of reconstruction, increased infrastructure efficiency, decreased economic costs of frequent repairs, and reduced environmental impacts of regular restorations [11].
@@
For these reasons, this study focuses on autogenous self-healing concrete, which more closely resembles traditional concrete and warrants investigation for real-world implementation. Autogenous healing concrete exhibits computational tractability compared to bacteria-based concrete due to its uniformity and predictability. All concrete types possess inherent autogenous self-healing abilities through continued hydration, carbonation, and calcium silicate hydrate (C-S-H) swelling. For this study, we define autogenous self-healing concrete as a cementitious mixture specifically modified to enhance self-healing capacity.
@@
Previous research has established a foundation for finite element analysis and physics-based modeling of autogenous self-healing concrete [15, 16, 17, 18]. However, several limitations persist across existing studies: (1) isotropic diffusion and healing assumptions that neglect anisotropic variations in concrete crystal matrices, (2) neglect of time-dependence and feedback mechanisms at each time step, (3) lack of experimental calibration or validation, and (4) simplified 2D domains that may not capture complex three-dimensional healing processes.
This paper presents a coupled simulation model that integrates multiple relevant physical and chemical processes previously overlooked. We developed two distinct model versions using the FEniCSx finite element software [19]: the Crack Diffusion Model (CDM) implementing standard diffusion across crack interfaces, and the Crack Membrane Model (CMM) implementing a binary gate that selectively allows diffusion only after reaching critical moisture levels. The models incorporate time-dependent feedback mechanisms, realistic diffusion coefficients, and damage evolution relating cement and water availability.
@@ 2.1 Crack Diffusion Model
where $U$ is the water concentration variable (dimensionless), $D$ is the diffusion coefficient ($\frac{m^{2}}{s}$), and $d$ is the damage variable (dimensionless). The damage variable represents the fraction of damaged area, where $d=0$ corresponds to intact concrete and $d=1$ to complete damage.
@@
where $\gamma$ is the characteristic smoothing length scale. The spatial domain and its boundary are denoted as $\Omega$ and $\partial\Omega$, respectively.
The CDM solves the coupled system of equations introduced above using the finite element method implemented in FEniCSx, with the workflow shown in Figure 1. The time-dependent solution accounts for feedback between moisture transport, damage evolution, and cement availability.
@@ 2.2 Crack Membrane Model
where $U_{\text{critical}}$ is the critical moisture threshold, $\Delta U_{\text{gate}}$ controls the transition sharpness, and $\epsilon_{\text{gate}}$ is a small positive constant that prevents numerical instabilities.
@@
This formulation suppresses water flow in cracked regions until the critical moisture threshold is reached, after which normal diffusion resumes. The CMM solves the coupled system of equations using the finite element method implemented in FEniCSx, with the workflow shown in Figure 2.
Figure 2: Crack Membrane Model Logic Workflow
@@ 2.3 Machine Learning Model
Simulations in this study require solving millions of partial differential equations on fine element meshes, demanding substantial computational resources. Machine learning models trained on physics-based simulation data can learn underlying physical relationships and provide predictions 330 to 2,960,000 times faster than standalone finite element analysis [20].
@@
The training of the ML models proceeded as follows. One million data points using various crack width indicator $\sigma$, cement availability $\gamma$, and time $t$ values were used to generate $H$, the healing variable that is equal to 0 when not fully healed and 1 when fully healed. Five standard classification models were trained to learn the binary classification function $H(\sigma,\gamma,t)$. Namely, they were the K-Nearest Neighbor (KNN), Logistic Regression, Naive Bayes, Support Vector Machine (SVM), and Neural Network Models. All of the models were trained out-of-the-box with the default parameters except for the neural network. Implementation utilized the TensorFlow and scikit-learn python libraries.
@@ 2.4 Simulation Cases and Preparation
This study used a 2D, $1\times 1$ cm unit square mesh discretized into 2,048 triangles of equal size. The initial condition was a pre-defined line crack passing through the center of the domain and having incline angle $\beta$ (Figure 4(a)). The impact of the crack angle on various properties and healing results was investigated. To initialize a continuous damage field, the baseline damage follows a Gaussian function as
@@
Figure 4: Initial Condition for the Damage Field: (a) Definition of inclining angle $\beta$, (b) Example Scalar Damage Field During Pre-Cracking ($\beta=0$, $\sigma^{2}=0.0005$)
@@
    What is the influence of crack angle and crack width on healing time?
@@
$D_{\text{intact}}$ Diffusivity of water in pure, healed concrete Constant: $10^{-8}\frac{cm^{2}}{s}$
$D_{\text{cracked}}$ Diffusivity of water in crack space, representing microcrack bridging, capillary bridges, partially healed CSH gel phases, and tortuous percolation Constant: $10^{-7}\frac{cm^{2}}{s}$ (except for the analysis of crack width on healing time)
@@ 3 Results and Analysis
To examine the complexity of healing behavior across different crack angles $\beta$, healing progression plots were generated for $\beta=0$, $\frac{\pi}{8}$, $\frac{\pi}{4}$, and $\frac{\pi}{2}$, as shown in Figure 7. Comparison between $\beta=0$ and $\beta=\frac{\pi}{8}$ reveals that the crack at angle $\beta=\frac{\pi}{8}$ initiated healing earlier than the vertical crack ($\beta=0$), as water reached the leftmost crack region sooner for $\beta=\frac{\pi}{8}$. However, the healing rate remained relatively slow because water had to continue saturating the domain until reaching the rightmost crack region of $\beta=\frac{\pi}{8}$. In contrast, once the vertical crack ($\beta=0$) began healing, it healed at a significantly faster rate because water could immediately begin traversing the entire crack width. Consequently, the vertical crack achieved complete healing in a shorter time due to reduced diffusion distance. For $\beta=0$ and $\frac{\pi}{8}$, the healing exhibited a symmetrical sigmoid shape. Comparison between the $\beta=\frac{\pi}{4}$ and $\beta=\frac{\pi}{2}$ cracks indicates that the horizontal crack ($\beta=\frac{\pi}{2}$) healed slightly faster than the $\beta=\frac{\pi}{4}$ crack. This observation is consistent with the Healing Time vs. Crack Angle plot in Figure 6, which demonstrates the influence of crack length on healing time when the crack mouth is in direct contact with the water source. Examination of the $\beta=\frac{\pi}{4}$ and $\frac{\pi}{2}$ graphs reveals rapid damage recovery at the beginning, then steadily increasing healing progress in the middle, and finally slowed healing progression as total healing approached 100%. Therefore, the 95% healed damage threshold was necessary; otherwise, healing times for full damage recovery would have been substantially longer, distorting the results. All graphs showed a 50% healing threshold at approximately $8\times 10^{5}$ seconds, representing the time required for water to reach the crack center. This observation further supports the conclusion that primarily diffusion determined healing time.
@@
The impact of the crack width on the healing time was investigated using the CDM. Two simulations were conducted with different parameter values. Within the range of $\sigma\in[0.005,0.040]$ cm, 37 values were chosen and simulated. For Simulation 1, its parameter values were $D_{\text{intact}}=10^{-8}$, $D_{\text{cracked}}=10^{-9}$, $\alpha$ = 0.01, $\beta$ = 0, and $\gamma$ = 0.0316. For Simulation 2, its parameter values were $D_{\text{intact}}=10^{-8}$, $D_{\text{cracked}}=10^{-7}$, $\alpha$ = 0.01, $\beta$ = 0, and $\gamma$ = 0.0316. The healing time versus crack width plots are shown in Figure 8.
@@
Figure 9: 3D Plot of Healing Time, Crack Width, and Smoothing Parameter
@@
The healing process is also of interest for practical use. Figure 10 illustrates the typical healing progression over time simulated using the CDM with parameter values $\alpha$ = 0.01, $\beta$ = 0, $\sigma$ = 0.0224, and $\gamma$ = 0.0316. The healing process demonstrated an initial slow phase followed by acceleration as water saturated the entire crack. The healing rate approached an asymptotic maximum until complete crack healing was achieved. This behavior reflects the rate-limiting nature of water diffusion in the healing process. As damage decreased, water diffusion efficiency increased, resulting in higher healing rates. Similar healing patterns were observed across all parameter configurations with the CDM.

### v2

@@ Abstract
A time-dependent modeling framework for autogenous self-healing concrete that couples moisture diffusion with damage evolution was developed. Water transport follows Fick’s second law with a damage-dependent diffusivity obtained by power-law interpolation between intact concrete and crack space. Healing reduces damage in proportion to the product of local moisture and a smoothed cement availability field computed via a novel Helmholtz filtering approach that models the spatial extent over which cement clinker can travel and form crystals. Two finite element variants were implemented in FEniCSx over time horizons up to $5\times 10^{6}$ seconds: a Crack Diffusion Model (CDM) with standard diffusion and a Crack Membrane Model (CMM) that introduces a novel threshold-based gating mechanism to control cross-crack water transport until a critical moisture threshold is reached. Key control parameters are the initial crack orientation and size, the diffusion coefficients of intact and cracked concrete, the healing rate constant, and the cement availability smoothing parameter. Simulations show that healing time varies non-monotonically with crack orientation, peaking near $45^{\circ}$ and $135^{\circ}$ and minimizing near $90^{\circ}$, consistent with diffusion distance to crack endpoints dominating the process. The dependence on crack width reverses with material parameters: healing time increases when $D_{\text{cracked}}<D_{\text{intact}}$ and decreases when $D_{\text{cracked}}>D_{\text{intact}}$. The CMM reproduces staged moisture penetration with delayed gate activation but lengthens total healing time, whereas the CDM is efficient for parametric sweeps. Machine learning regression models were trained on finite element simulation data to predict healing times $H(\sigma,\gamma,\beta)$ with high accuracy ($R^{2}>0.999$), dramatically reducing computational time. SHAP analysis demonstrated that crack orientation influenced healing time the most, followed by crack width and cement availability smoothing. Although experimental calibration is still required, the framework provides a versatile tool for guiding future laboratory studies and implementations of self-healing concrete.
Keywords: self-healing concrete; finite element analysis; machine learning; autogenous healing; diffusion; damage evolution; FEniCSx.
@@ 1 Introduction
1.1 Background
Concrete is the second most widely used material globally, surpassed only by water, with over 30 billion metric tons utilized worldwide annually [1]. This critical construction material has been employed for over 8,000 years, dating back to concrete-like structures built by Nabataean traders in modern-day Syria [2], and continues to be essential for housing, transportation, and water infrastructure today. The global concrete market is projected to reach \$821.6 billion by 2026 [3], driven by growing investment in construction projects worldwide.
@@ 1.1 Background
The infrastructure crisis compounds these environmental challenges. Most countries completed their major infrastructure projects during the late 1900s, including the US Interstate Highway System, European motorways, Japan’s Shinkansen, and China’s expressway networks. Over half a century later, these aging structures face increasing failure risks, with concrete failure probability approaching 70% after 50 years and becoming nearly inevitable by the century’s end [7]. Currently, over 20,000 concrete bridges in the United States are classified as structurally deficient [8], requiring frequent maintenance and posing safety risks.
Crack formation represents an inherent characteristic of concrete structures, arising from multiple mechanisms: structural loading that exceeds concrete’s limited tensile strength, volumetric changes associated with shrinkage and expansion, and environmental degradation processes including corrosion and freeze-thaw cycles [9, 10]. The presence of cracks significantly compromises concrete durability by providing pathways for liquids and gases to penetrate to the steel reinforcement, thereby accelerating corrosion processes. The global cost of corrosion in reinforced concrete infrastructure is estimated at \$2.5 trillion annually, representing approximately 3.4% of global GDP [11].
Self-healing concrete emerges as a promising solution to address these challenges. This innovative material repairs its cracks automatically without human intervention, typically through chemical or biological processes [12]. Self-healing concrete offers reduced maintenance requirements over the infrastructure lifetime, diminished social costs of reconstruction, increased infrastructure efficiency, decreased economic costs of frequent repairs, and reduced environmental impacts of regular restorations [13].
@@
For these reasons, this study focuses on autogenous self-healing concrete, which more closely resembles traditional concrete and offers greater potential for real-world implementation. Autogenous self-healing concrete exhibits superior computational tractability compared to bacteria-based systems due to its uniformity and predictability. All concrete types possess inherent autogenous self-healing capabilities through continued hydration, carbonation, and calcium silicate hydrate (C-S-H) swelling, although the healed material generally exhibits inferior properties compared to the original material [9, 10]. For most multi-mineral materials, optimal self-healing performance is achieved through a combination of silica-based swelling and crystalline components, enabling crack repair up to 0.6 millimeters in width [10, 16]. In this study, autogenous self-healing concrete is defined as a cementitious mixture specifically modified to enhance self-healing capacity.
@@
1.2 Framework and Novelty of the Study
We developed two distinct model versions using the FEniCSx finite element software [18]: the Crack Diffusion Model (CDM) implementing standard diffusion across crack interfaces, and the Crack Membrane Model (CMM) implementing a binary gate that selectively allows diffusion only after reaching critical moisture levels. The models incorporate time-dependent feedback mechanisms, realistic diffusion coefficients, and damage evolution that couples cement and water availability.
Previous research has established a foundation for finite element analysis and physics-based modeling of autogenous self-healing concrete [19, 20, 21, 22]. Building upon this foundation, the present work addresses several areas where enhanced modeling capabilities could provide more accurate predictions.
First, while some studies have employed thermodynamic approaches where healing is governed by strain and energy equivalence [21, 23], autogenous self-healing concrete is driven by the hydration of cement clinker and the diffusion of water into the crack space. A diffusion-based approach captures the chemical processes that form healing crystals and accounts for the limiting factors of water and cement clinker availability. The models implement damage-dependent anisotropic diffusion, where the diffusion coefficient varies spatially based on the local damage field. Additionally, while some previous studies have assumed continuous water transport across crack interfaces [19], the CMM incorporates a gating mechanism that requires sufficient moisture buildup before allowing cross-crack transport, representing the physical process of pressure-based water penetration.
The healing reactions depend on both the availability of water and cement clinker. While some studies have assumed uniform healing rates or simplified kinetics [24, 25, 26], the CDM and CMM incorporate a kinetics relationship that captures both water and cement clinker availability. The cement availability smoothing parameter accounts for both the spatial distance over which cement can be transported and its non-uniform, damage-dependent availability.
The models fully couple diffusion and reaction processes within a single time-dependent framework, whereas previous studies have typically implemented these processes independently [27]. By updating each equation at each time step, the models capture feedback mechanisms that combine various physical processes into a unified simulation. This fully coupled approach represents the interdependent nature of moisture transport, cement availability, and healing kinetics. To address the computational cost of the coupled models, machine learning is used to predict healing times efficiently [28]. Explainable AI (XAI) is used to further interpret the models’ predictions.
@@ 2.1 Crack Diffusion Model
where $U$ is the water concentration variable (dimensionless), $D(d)$ is the diffusion coefficient ($\frac{\text{cm}^{2}}{\text{s}}$) as a function of $d$, and $d$ is the damage variable (dimensionless). The damage variable represents the fraction of damaged area, where $d=0$ corresponds to intact concrete and $d=1$ to complete damage.
@@
where the spatial domain and its boundary are denoted as $\Omega$ and $\partial\Omega$, respectively, and $\gamma$ is the characteristic smoothing length scale. The smoothing parameter describes the distance over which cement clinker can be transported by water to form healing crystals.
The CDM solves the coupled system of equations introduced above using the latest version of FEniCSx in Python, with the workflow shown in Figure 1. The time-dependent solution accounts for feedback between moisture transport, damage evolution, and cement availability.
@@ 2.2 Crack Membrane Model
where $U_{\text{critical}}$ is the critical moisture threshold, $\Delta U_{\text{gate}}$ controls the transition sharpness, and $\epsilon_{\text{gate}}$ is a small positive constant that prevents numerical instabilities. When $U>>U_{\text{critical}}$, $G(U)=1$, allowing for normal diffusion. When $U<<U_{\text{critical}}$, $G(U)=0$, preventing water transport.
@@
This formulation suppresses water flow in cracked regions until the critical moisture threshold and water pressure are reached, after which normal diffusion across the crack resumes. The CMM also solves the coupled system of equations using the latest version of FEniCSx in Python, with the workflow shown in Figure 2.
Figure 2: Crack Membrane Model Logic Workflow for Water Transport Across Cracks
@@ 2.3 Machine Learning Model
Simulations in this study require solving millions of partial differential equations on fine meshes, demanding substantial computational resources. Machine learning models trained on physics-based simulation data can learn underlying physical relationships and provide predictions 330 to 2,960,000 times faster than standalone finite element analysis [29].
@@
The training of the ML models proceeded as follows. A dataset of 1100 simulation samples was generated from the CDM by systematically varying the crack width indicator $\sigma$, cement availability smoothing length $\gamma$, and crack orientation $\beta$ over the parameter ranges detailed in Table 1. All parameters were normalized to zero mean and unit variance using z-score normalization. For each parameter combination, the healing time $H$ was computed as the time required to heal 95% of the damage. Other parameters were kept constant. Specifically, their values were $\alpha$ = 0.01, $D_{\text{intact}}$ = $10^{-8}\frac{\text{cm}^{2}}{\text{s}}$, and $D_{\text{cracked}}$ = $10^{-7}\frac{\text{cm}^{2}}{\text{s}}$. The dataset was partitioned into training (60%), validation (20%), and test (20%) subsets for model development and evaluation. Five standard regression models were trained to learn the regression function $H(\sigma,\gamma,\beta)$. Namely, they were the Gaussian Process, Ridge Regression, Random Forest, eXtreme Gradient Boosting (XGBoost), and Neural Network Models. All of the models were trained out-of-the-box with the default parameters except for the neural network. For the neural network, a 3-32-32-1 architecture with ReLU activation, dropout=0.2, learning rate=0.001, batch size=32, and 100 epochs was used. To interpret model predictions and quantify feature importance, SHapley Additive exPlanations (SHAP) values were computed for each input feature. The feature contributions are visualized using both bar charts and summary plots. Implementation utilized the latest releases of the TensorFlow, scikit-learn, XGBoost, and SHAP libraries in Python.
Parameter Lower Bound Upper Bound Step Size Number of Steps
$\sigma$ $005\text{ cm} 0.0$$5\text{ cm} 0.0$$005\text{ cm}$ $10$
$\gamma$ $005\text{ cm} 0.0$$5\text{ cm} 0.0$$005\text{ cm}$ $10$
$\beta$ 0 $\frac{\pi}{2}$ $\frac{\pi}{20}$ $11$
@@ 2.4 Simulation Cases and Preparation
This study used a 2D, $1\times 1$ cm unit square mesh discretized into 2,048 triangles of equal size. The initial condition was a pre-defined line crack passing through the center of the domain with inclination angle $\beta$ (Figure 4(a)). The impact of the crack angle on various properties and healing results was investigated. To initialize a continuous damage field, the baseline damage follows a Gaussian function as
@@
Figure 4: Initial Condition for the Damage Field: (a) Definition of Inclination Angle $\beta$, (b) Example Scalar Damage Field During Pre-Cracking ($\beta=0$, $\sigma^{2}=0.0005$)
@@
    What is the influence of crack angle, crack width, and cement availability on healing time?
@@
$D_{\text{intact}}$ Diffusivity of water in pure, healed concrete Constant: $10^{-8}\frac{\text{cm}^{2}}{\text{s}}$
$D_{\text{cracked}}$ Diffusivity of water in crack space, representing microcrack bridging, capillary bridges, partially healed C-S-H gel phases, and tortuous percolation Constant: $10^{-7}\frac{\text{cm}^{2}}{\text{s}}$ (except for the analysis of crack width on healing time)
@@
Animation videos were generated in Pyvista, with CDM damage evolution for various crack angles shown in Figure 5 and CMM damage and saturation evolutions shown in Figure 6.
Figure 5: Video 1. CDM Damage Evolution Animations for $\beta=0,\frac{\pi}{8},\frac{\pi}{2}$
Figure 6: Video 2. CMM Damage and Saturation Evolution Animations
@@ 3 Results and Analysis
To examine the complexity of healing behavior across different crack angles $\beta$, healing progression plots were generated for $\beta=0$, $\frac{\pi}{8}$, $\frac{\pi}{4}$, and $\frac{\pi}{2}$, as shown in Figure 9. Comparison between $\beta=0$ and $\beta=\frac{\pi}{8}$ reveals that the crack at angle $\beta=\frac{\pi}{8}$ initiated healing earlier than the vertical crack ($\beta=0$), as water reached the leftmost crack region sooner for $\beta=\frac{\pi}{8}$. However, the healing rate remained relatively slow because water had to continue saturating the domain until reaching the rightmost crack region of $\beta=\frac{\pi}{8}$. In contrast, once the vertical crack ($\beta=0$) began healing, it healed at a significantly faster rate because water could immediately begin traversing the entire crack width. Consequently, the vertical crack achieved complete healing in a shorter time due to reduced diffusion distance. For $\beta=0$ and $\frac{\pi}{8}$, the healing exhibited a symmetrical sigmoid shape. Comparison between the $\beta=\frac{\pi}{4}$ and $\beta=\frac{\pi}{2}$ cracks indicates that the horizontal crack ($\beta=\frac{\pi}{2}$) healed slightly faster than the $\beta=\frac{\pi}{4}$ crack. This observation is consistent with the Healing Time vs. Crack Angle plot in Figure 8, which demonstrates the influence of crack length on healing time when the crack mouth is in direct contact with the water source. Examination of the $\beta=\frac{\pi}{4}$ and $\frac{\pi}{2}$ graphs reveals rapid damage recovery at the beginning, then steadily increasing healing progress in the middle, and finally slowed healing progression as total healing approached 100%. Therefore, the 95% healed damage threshold was necessary; otherwise, healing times for full damage recovery would have been substantially longer, distorting the results. All graphs showed a 50% healing threshold at approximately $8\times 10^{5}$ seconds, representing the time required for water to reach the crack center. This observation further supports the conclusion that healing time is primarily determined by diffusion processes.
Crack orientation has not been analyzed as a factor in diffusion-based autogenous healing models. Previous studies typically assume cracks are perpendicular to the diffusion direction or that the crack mouth is in direct contact with the water source. This work examines healing behavior across a range of crack orientations.
@@
The impact of the crack width on the healing time was investigated using the CDM. Two simulations were conducted with different parameter values. Within the range of $\sigma\in[0.005,0.040]$ cm, 37 values were chosen and simulated. For Simulation 1, its parameter values were $D_{\text{intact}}=10^{-8}$ cm2/s, $D_{\text{cracked}}=10^{-9}$ cm2/s, $\alpha$ = 0.01, $\beta$ = 0, and $\gamma$ = 0.0316. For Simulation 2, its parameter values were $D_{\text{intact}}=10^{-8}$ cm2/s, $D_{\text{cracked}}=10^{-7}$ cm2/s, $\alpha$ = 0.01, $\beta$ = 0, and $\gamma$ = 0.0316. The healing time versus crack width plots are shown in Figure 10.
@@
Figure 11: 3D Plot of Healing Time $t$, Crack Width $\sigma$, and Smoothing Parameter $\gamma$
@@
Previous studies have shown that the healing process is highly dependent on crack width [30]. For wider cracks, healing has been shown to be diffusion-limited [20], consistent with our finding that increasing crack width decreases the diffusion distance and time to the crack boundary. The cement smoothing parameter introduced here has not been implemented in previous models.
@@
The healing process is also of interest for practical use. Figure 12 illustrates the typical healing progression over time simulated using the CDM with parameter values $\alpha$ = 0.01, $\beta$ = 0, $\sigma$ = 0.0224, and $\gamma$ = 0.0316. The healing process demonstrated an initial slow phase followed by acceleration as water saturated the entire crack. The healing rate approached an asymptotic maximum until complete crack healing was achieved. This behavior reflects the rate-limiting nature of water diffusion in the healing process. As damage decreased, water diffusion efficiency increased, resulting in higher healing rates. Similar healing patterns were observed across all parameter configurations with the CDM. Previous studies have demonstrated that crack closure does not occur under low-humidity conditions. In contrast, under wet-dry cycling conditions and sustained immersion, healing efficiency exhibits a progressive increase over time until reaching a maximum value [31]. These observations are consistent with our model predictions, which indicate that elevated water saturation levels correspond to enhanced healing rates.